# The Self-Organized Criticality Paradigm in Economics & Finance

### v1

@@ Abstract
“Self-Organised Criticality” (SOC) is the mechanism by which complex systems spontaneously settle close to a critical point, at the edge between stability and chaos, and characterized by fat-tailed fluctuations and long-memory correlations. Such a scenario may explain why insignificant perturbations can generate large disruptions, through the propagation of “avalanches” across the system. In this short review, we discuss how SOC could offer a plausible solution to the excess volatility puzzle in financial markets and the analogue “small shocks, large business cycle puzzle” for the economy at large, as initially surmised by Per Bak et al. in 1993 [1]. We argue that in general the quest for efficiency and the necessity of resilience may be mutually incompatible and require specific policy considerations.
@@ 1 Introduction
Many systems made up of a large number of interacting items appear to be “marginally stable”, i.e. close to an incipient instability, or tipping point. The scenario according to which these systems are spontaneously driven towards such a fragile state is called “self-organised criticality” (SOC), and was proposed by Per Bak [2] as a generic mechanism that allows one to explain large endogenous fluctuations in complex systems, ranging from natural systems (avalanches, earthquakes, floods, turbulence…) [3, 4, 5, 6, 7, 8], ecological systems (mass extinctions) [9, 10], biological neural networks (epilepsy) [11, 12, 13, 14], bird flocks and fish schools (collective motion) [15], socio-technical systems (black-outs, traffic jams, failure cascades,…) [16, 17, 18], etc., etc.
@@
Of course, as the French saying goes, comparaison n’est pas raison. It may well be that all these analogies are misleading, and that the true underlying mechanisms leading to excess volatility should be found elsewhere. For one thing, some authors would still argue that volatility is not excessive at all, but merely reflects the existence of genuine exogenous shocks of large amplitude, perhaps invisible to the economist, but adequately processed by smart market participants and economic agents. Needless to say, such an “invisible hand” explanation often sounds preposterous, but reflects that it is hard to accept (and quite disturbing indeed) that large events can occur without “large causes” – whereas it is the major epistemological lesson of complexity science in general, and SOC in particular, that tiny perturbations can induce full crises. A complex system can actually be defined as a system where small perturbations can – but not necessarily do – trigger incommensurate effects, see e.g. [34, 35, 6, 36].
Other explanations have been proposed, more plausible than phantom shocks. For example, the “Granularity Hypothesis” (GH) of X. Gabaix [37] is an attempt to rationalize the fact that idiosyncratic shocks may survive at the aggregate level, because some firms have a disproportionate size and cannot be averaged out (see also [38]). This is the case when firm sizes $S$ are power-law distributed, with a tail decaying slow enough, for example as $S^{-2}$ as is indeed the case empirically [39]. Similarly, according to [40], large price jumps in financial markets would result from orders sent by large investors or funds, the size of which is also known to be power-law distributed [41].
@@
Yet another mechanism, that we will discuss further in sections 3.4 and 4.3, is that the equilibrium state of the system is in fact dynamically unstable, leading to a quasi-periodic or chaotic evolution even in the absence of any exogenous shock. Excess volatility is then purely self-induced. In this case, one does not expect to see generic power-law statistics or long-memory effects when the system is left on its own device. Interestingly, however, attempts to stabilize the system (through e.g. regulation, or monetary policy, etc.) might yet again drive the system close to a critical point, reinstalling, in a sense, the SOC scenario.
Finally, let us briefly mention a related but different concept: Self-Organized Bistability [42], which has been argued to be relevant in different situations. Such a scenario might be also interesting to consider in the context of economic or financial systems, see e.g. [43] for similar ideas.
The outline of this paper is as follows. We first explain in section 2 what is special about critical (marginally unstable) points. We then motivate the concept of Self-Organized Criticality in section 3 and give a few examples where such a concept naturally applies. We next turn to possible applications of SOC in economical (section 4) and financial (section 5) contexts. Finally, we discuss in the conclusion section 6 the policy implications of the possible fragility of socio-economic systems. We argue in particular that efficiency and resilience are often incompatible, and that operators, regulators and policy makers should take stock of the unintended consequences that can appear when resilience is not explicitly included in the welfare function.
@@ 2.2 Critical Branching
When $R_{0}<1$, a first unstable grain dislodges on average a finite number of other grains. More precisely, the average size of avalanches is given by $(1-R_{0})^{-1}$. The probability of very large avalanches is exponentially small. However, as $R_{0}$ approaches $1$, one single grain can lead to large disruptions. The probability of triggering an avalanche involving $S$ grains is given by [44]
@@
In other words, when $R_{0}=1$ the distribution of avalanche sizes is a scale-free, power-law distribution $S^{-3/2}$, with infinite mean. [^2] Note that when the distribution of off-springs $\rho(n;R_{0})$ decays as $n^{-1-\alpha}$ with $21<\alpha<$ 2, the critical avalanche size distribution has a power-law tail given by $S^{-1-1/\alpha}$. Still, the probability $\phi$ that the avalanche continues forever is zero.
@@ 3.1 Motivation
The seminal idea of Per Bak is to think of model parameters themselves as dynamical variables, in such a way that the system spontaneously evolves towards the critical point, or at least visits its neighbourhood frequently enough (for reviews, see [2, 45, 8]). Take for example the case of highly transmissible virus, with a natural value of $R_{0}$ above unity. As the disease spreads, containment measures are adopted (masks, social distancing, etc.) and the risk of getting sick is taken seriously by the population. This leads to a significant reduction of $R_{0}$, perhaps even below unity. Now, the epidemics slows down and people become more complacent, which causes $R_{0}$ to rise again. It is not absurd to imagine that the process will indeed self-organize around a point where the spread of the disease is curbed at a minimal amount of individual constraints. This is achieved precisely at the critical value $R_{0}=1$.
@@ 3.2 Sweeping through an instability
Consider now the sand pile example initially put forth by Per Bak and collaborators [3]. Start with a flat layer of sand and slowly add grains from a point source above it. A conical pile will form, with a progressively steeper slope. As the slope increases, so does the probability that a rolling grain destabilizes more grains, i.e. $R_{0}$ increases. As long as $R_{0}<1$, the resulting avalanches are small and do not change the overall shape of the pile. But as soon as $R_{0}>1$, there is a possibility of a landslide that collapses the pile and reduces its slope. The dynamical evolution of $R_{0}$ can then be described as [^3] Throughout the paper we use the standard notation $(x)^{+}=\max(x,0)$.
@@
where $P(R_{0},t)$ is the probability to find a slope characterized by a certain value of $R_{0}$, $\mu$ describes the rate at which $R_{0}$ increases due to the addition of new grains, and $\gamma$ the rate at which single grains start rolling down the slope, and $(R_{0}-1)^{+}$ the probability that a single grain triggers a landslide. For simplicity, one assumes that after the landslide, the pile restarts at $R_{0}=0$.
@@
with $Z=1+\sqrt{{\pi\mu}/{2\gamma}}$ such that $Q_{\text{st.}}(R_{0})$ is normalized to one. This equation shows that there is a finite probability for the slope to be exactly critical. This is enough to ensure that one will observe a power-law distribution of avalanche sizes [46], which can be computed from
@@
where $P(S)$ is given by Eq. (4). It is not very difficult to check that $\mathcal{P}(S)$ behaves at large $S$ as $S^{-2}$ without any exponential truncation term as in Eq. 4, i.e. as a pure power-law. Note however that $\mathcal{P}(S)$ only captures avalanches of finite size (i.e. much smaller than the total size of the system). One should also consider system-wide avalanches which relax the system all the way down to $R_{0}=0$. These contribute to an additional hump for very large $S$ of the order of the system size – events that were called “Dragon Kings” in [47] (see Fig. 1).
@@
The above example illustrates one possible scenario through which intermittent dynamics, with a power-law distribution of event sizes, can occur. The critical point is regularly visited by the natural dynamics of the system, as it travels from the stable phase to the unstable phase, the instability driving the system back to the stable phase. As discussed below, such a scenario is also reminiscent of the Minsky cycle in financial markets.
@@ 3.3 Dynamical convergence towards criticality
A different scenario is when the natural dynamics stops precisely at the critical point. For example, gradient descent applied to complex Constrained Satisfaction Problems generically stops when a local minimum of the loss function is reached. In many cases however, such minima are marginally stable, that is, the Hessian of the loss function has all its eigenvalues non negative, but the smallest one is very close to zero, see e.g. [48, 49, 50]. This is precisely the situation described by Eq. (3), where matrix $\mathbb{K}$ is the Hessian matrix. Therefore, small shocks get amplified by the proximity of an instability and small static perturbations can disproportionately modify the equilibrium state. In fact, such perturbations can even change the sign of the smallest eigenvalue, forcing the system to rearrange and find another (marginally stable) equilibrium, often very different from the original (unperturbed) one.
As an illustration, let us consider the so-called Generalized Lotka-Volterra model, describing the population dynamics of many interacting species [51]. We denote $x_{i}(t)$ the number of living individuals of species $i=1,\ldots,N$ and write
@@
where $\mu_{i}>0$ is the “fitness” of specie $i$ (i.e. its growth rate in the absence of interaction with other species, including itself) and the matrix $\mathbb{A}$ captures beneficial ($\mathbb{A}_{ij}>0$) or detrimental ($\mathbb{A}_{ij}<0$) interactions with other species. (For example, $\mathbb{A}_{ij}\mathbb{A}_{ji}<0$ describes a predator-prey situation).
The ecological equilibria corresponding to Eq. (7) must be such that either $x_{i}^{\star}=0$ (extinct specie) or $\mu_{i}+\sum_{j=1}^{N}\mathbb{A}_{ij}x_{j}^{\star}=0$. Naively solving the second equation yields, in a vectorial notation
@@
However, the resulting vector $\vec{x}^{\star}$ will in general contain negative entries, which of course does not make sense since the number of living individuals must be positive or zero. Feasible equilibria are thus such that only a subset of species survives, such that Eq. (8) restricted to that subset returns a vector $\vec{x}^{\star}$ with only positive entries. As first anticipated by the late Sir Robert May, not all ecologies can be stable [9].
@@
has a diverging eigenvalue $\lambda^{\star}$ in the large $N$ limit [52]. This means that any small change in the fitness of one specie can have dramatic consequences on the whole system – in the present case, mass extinctions [10]. Similarly, the stability matrix $\mathbb{K}$, given by (cf. Eq. (3))
@@ 3.4 Stabilizing an unstable equilibrium
In other words, the better one is able to stabilize the system, the more difficult it becomes to predict its future evolution! This in turn generates occasional large errors, leading to large swings that are later corrected as the control policy becomes efficient again [61]. This rather beautiful scenario is quite universal – for example heart-beat regulation is a result of two conflicting forces: excitation (sympathetic nervous system) and inhibition (parasympathetic nervous system), or brain activity regulation [65]. Such ideas are probably also relevant in economics and finance (see section 5.2, 5.4 below). In fact, the title of the present subsection is tantalizingly close to that of Minsky’s seminal book, “Stabilizing an Unstable Economy” [66].
@@ 4 Self-Organized Criticality in Economics
The aim of the present section is to revisit the issue and discuss more recent proposals that lead to self-organized criticality in economic systems, based on arguably less artificial assumptions. The hope is that putting the SOC scenario on firmer theoretical ground could revive the interest of economists. It is indeed fair to say that despite the high profile of its authors, the 1993 paper of Bak et al. [1] did not reach the level of attention that (in my view) it deserved. [^4] At the time of writing, the joint number of citations of Bak et al. 1993 [1] and Scheinkman-Woodford 1994 [21] is $\sim 800$ according to Google Scholar, compared for example to 2,050 for Gabaix’ 2011“granularity” paper [37], or 2,700 for Acemoglu, Carvalho et al. 2012 paper on the network origin of aggregate fluctuations [67].
@@ 4.2 The ecology of production networks (statics)
Expanding Eq. (11) to first order in $\zeta$, one notices that since $\sum_{j>0}a_{ij}<1$, the Perron-Frobenius theorem ensures that $q\to\infty$ Cobb-Douglas networked economies (such as those considered [67] always have a feasible equilibrium where all firms can survive, for any network and any productivities. Therefore, the type of crisis that takes place for $q<+\infty$, where bankruptcies occur, has no counterpart in a Cobb-Douglas economy.

### v2

@@ Abstract
“Self-Organised Criticality” (SOC) is the mechanism by which complex systems are spontaneously driven towards, or even across, a critical point at the edge between stability and chaos. These special points are characterized by fat-tailed fluctuations and long-memory correlations. Such a scenario can explain why insignificant perturbations may generate large disruptions, through the propagation of “avalanches” across the system. In this short review, we discuss how SOC could offer a plausible solution to the excess volatility puzzle in financial markets and the analogue “small shocks, large business cycle puzzle” for the economy at large, as initially surmised by Per Bak et al. in 1993 [1] and, in a different language, by Hyman Minsky. We argue that in general the quest for efficiency and the necessity of resilience may be mutually incompatible and require specific policy considerations.
@@ 1 Introduction
Many systems made up of a large number of interacting items appear to be “marginally stable”, i.e. close to an incipient instability, or tipping point. The scenario according to which these systems are spontaneously driven towards such a fragile state is called “self-organised criticality” (SOC), and was proposed by Per Bak [2] as a generic mechanism that allows one to explain large endogenous fluctuations in complex systems, ranging from natural systems (avalanches, earthquakes, floods, solar flares, turbulence…) [3, 4, 5, 6, 7, 8], ecological systems (mass extinctions) [9, 10], biological neural networks (epilepsy) [11, 12, 13, 14], bird flocks and fish schools (collective motion) [15], socio-technical systems (black-outs, traffic jams, failure cascades,…) [16, 17, 18], etc., etc.
@@
The aim of this (short) review is to discuss whether such a paradigm makes sense in the context of economic systems and financial markets. Indeed, as the French saying goes, comparaison n’est pas raison. It may well be that all these analogies are misleading, and that the true underlying mechanisms leading to excess volatility should be found elsewhere. For one thing, some authors would still argue that volatility is not excessive at all, but merely reflects the existence of genuine exogenous shocks of large amplitude, perhaps invisible to the economist, but adequately processed by smart market participants and economic agents. Needless to say, such an “invisible hand” explanation often sounds preposterous, but reflects that it is hard to accept (and quite disturbing indeed) that large events can occur without “large causes” – whereas it is the major epistemological lesson of complexity science in general, and SOC in particular, that tiny perturbations can induce full crises. A complex system can actually be defined as a system where small perturbations can – but not necessarily do – trigger incommensurate effects, see e.g. [34, 35, 6, 36].
Other explanations have been proposed, which are more plausible than phantom shocks. For example, the “Granularity Hypothesis” (GH) of X. Gabaix [37] is an attempt to rationalize the fact that idiosyncratic shocks may survive at the aggregate level, because some firms have a disproportionate size and cannot be averaged out (see also [38]). This is the case when firm sizes $S$ are power-law distributed, with a p.d.f. tail decaying slow enough, for example as $S^{-2}$ as is indeed the case empirically [39]. Similarly, according to [40], large price jumps in financial markets would result from orders sent by large investors or funds, the size of which is also known to be heavy-tailed [41] (but probably not power-law tailed, see [42]).
@@
Yet another mechanism, that we will discuss further in sections 3.4 and 4.3, is that the equilibrium state of the system is in fact dynamically unstable, leading to a quasi-periodic or chaotic evolution even in the absence of any exogenous shock. Excess volatility is then purely self-induced. As illustrated by the model of section 4.3, one does not necessarily expect to see power-law statistics or long-memory effects when such systems are left on their own device. Interestingly, however, attempts to stabilize the system (through e.g. regulation, monetary policy, learning, etc.) might yet again drive the system close to a critical point, reinstalling, in a sense, the SOC scenario.
Finally, let us mention two related but different concepts that we will not develop further here: Highly Optimized Tolerance (HOT) [43] and Self-Organized Bistability (SOB) [44], which have been argued to be relevant in different situations. Such scenarios might be also interesting to consider in the context of economic or financial systems, see e.g. [45] for similar ideas, and section 6.
The outline of this paper is as follows. We first explain in section 2 what is special about critical (marginally unstable) points. We then motivate the concept of Self-Organized Criticality in section 3 and give a few examples where such a concept naturally applies. We next turn to possible applications of SOC in economical (section 4) and financial (section 5) contexts. Finally, we discuss in the conclusion section 6 the policy implications of the possible fragility of socio-economic systems. We argue in particular that efficiency and resilience are often incompatible, and that operators, regulators and policy makers should take stock of the unintended consequences that can appear when resilience (i.e. tolerance to tail events) is not explicitly included in the welfare function.
@@ 2.2 Critical Branching
When $R_{0}<1$, a first unstable grain dislodges on average a finite number of other grains. More precisely, the average size of avalanches is given by $(1-R_{0})^{-1}$. The probability of very large avalanches is exponentially small. However, as $R_{0}$ approaches $1$, one single grain can lead to large disruptions. The probability of triggering an avalanche involving $S\gg 1$ grains is given by [46]
@@
In other words, when $R_{0}=1$ the distribution of avalanche sizes is a scale-free, power-law distribution $S^{-3/2}$, with infinite mean. [^2] Note that when the distribution of off-springs $\rho(n;R_{0})$ decays as $n^{-1-\alpha}$ with $21<\alpha<$ 2, the critical avalanche size distribution has a power-law tail given by $S^{-1-1/\alpha}$. Note that for $\alpha=2$ one recovers the classic $S^{-3/2}$ behaviour. Still, the probability $\phi$ that the avalanche continues forever is zero.
@@ 3.1 Motivation
The seminal idea of Per Bak is to think of model parameters themselves as dynamical variables, in such a way that the system spontaneously evolves towards the critical point, or at least visits its neighbourhood frequently enough (for a precursor of this general idea, see section 8 of Ref. [47], and for reviews, see [2, 48, 8]). Take for example the case of highly transmissible virus, with a natural value of $R_{0}$ above unity. As the disease spreads, containment measures are adopted (masks, social distancing, etc.) and the risk of getting sick is taken seriously by the population. This leads to a significant reduction of $R_{0}$, perhaps even below unity. Now, the epidemics slows down and people become more complacent, which causes $R_{0}$ to rise again. It is not absurd to imagine that the process will indeed self-organize around a point where the spread of the disease is curbed at a minimal amount of individual constraints. This is achieved precisely at the critical value $R_{0}=1$.
@@ 3.2 Sweeping through an instability
Consider now the sand pile example initially put forth by Per Bak and collaborators [3]. Start with a flat layer of sand and slowly add grains from a point source above it. A conical pile will form, with a progressively steeper slope. As the slope increases, so does the probability $R_{0}$ that a rolling grain destabilizes more grains increases. We will thus describe the dynamics of the pile in terms of $R_{0}$ rather than its slope.
As long as $R_{0}<1$, the resulting avalanches are small and do not change the overall shape of the pile. But as soon as $R_{0}>1$, there is a possibility of a landslide that collapses the pile and reduces its slope. The dynamical evolution of the probability to find a pile with “slope” $R_{0}$ is then described as [^3] Throughout the paper we use the standard notation $(x)^{+}=\max(x,0)$.
@@
where $P(R_{0},t)$ is the probability to find a slope characterized by a certain value of $R_{0}$, $\mu$ describes the rate at which $R_{0}$ increases due to the addition of new grains, and $\gamma$ the rate at which single grains start rolling down the slope, and $(R_{0}-1)^{+}$ the probability that a single grain does trigger a landslide. For simplicity, one assumes that after the landslide, the pile restarts at $R_{0}=0$.
@@
with $Z=1+\sqrt{{\pi\mu}/{2\gamma}}$ such that $Q_{\text{st.}}(R_{0})$ is normalized to one. This equation shows that there is a finite probability density for the slope to be exactly critical. This is enough to ensure that one will observe a power-law distribution of avalanche sizes [49], which can be computed from
@@
where $P(S)$ is given by Eq. (4). It is not very difficult to check that $\mathcal{P}(S)$ behaves at large $S$ as $S^{-2}$ without the exponential truncation term present in Eq. (4), i.e. as a pure power-law with exponent $-2$ instead of $-3/2$. Note however that $\mathcal{P}(S)$ only captures avalanches of finite size (i.e. much smaller than the total size of the system). One should also consider system-wide avalanches which relax the system all the way down to $R_{0}=0$. These contribute to an additional hump for very large $S$ of the order of the system size – events that were called “Dragon Kings” in [50] (see Fig. 1).
@@
The above example illustrates one possible scenario through which intermittent dynamics, with a power-law distribution of event sizes, can occur (see also [47] for an early discussion). The critical point is regularly visited by the natural dynamics of the system, as it travels from the stable phase to the unstable phase, the instability driving the system back to the stable phase. As discussed below, such a scenario is also reminiscent of the Minsky cycle in financial markets.
@@ 3.3 Dynamical convergence towards criticality
A different scenario is when the natural dynamics stops precisely at the critical point. For example, gradient descent applied to complex Constrained Satisfaction Problems generically stops when a local minimum of the loss function is reached. In many cases however, such minima are marginally stable, that is, the Hessian of the loss function has all its eigenvalues non negative, but the smallest one is very close to zero, see e.g. [51, 52, 53]. This is precisely the situation described by Eq. (3), where stability matrix $\mathbb{K}$ is the Hessian matrix. Therefore, small shocks get amplified by the proximity of an instability and small static perturbations can disproportionately modify the equilibrium state (see also [43] for related ideas). In fact, such perturbations can even change the sign of the smallest eigenvalue, forcing the system to rearrange and find another (marginally stable) equilibrium, often very different from the original (unperturbed) one.
As an illustration, let us consider the so-called Generalized Lotka-Volterra model, describing the population dynamics of many interacting species [54]. We denote $x_{i}(t)$ the number of living individuals of species $i=1,\ldots,N$ and assume that these numbers evolve as
@@
where $\mu_{i}>0$ is the “fitness” of species $i$ (i.e. its growth rate in the absence of interaction with other species, including itself) and the matrix $\mathbb{A}$ captures beneficial ($\mathbb{A}_{ij}>0$) or detrimental ($\mathbb{A}_{ij}<0$) interactions with other species. (For example, $\mathbb{A}_{ij}\mathbb{A}_{ji}<0$ describes a predator-prey situation).
The ecological equilibria corresponding to Eq. (7) must be such that either $x_{i}^{\star}=0$ (extinct specie) or $\mu_{i}+\sum_{j=1}^{N}\mathbb{A}_{ij}x_{j}^{\star}=0$. Naively solving the second equation yields, in a vector notation
@@
However, the resulting vector $\vec{x}^{\star}$ will in general contain negative entries, which of course does not make sense since the number of living individuals must be positive or zero. Feasible equilibria are thus such that only a subset of species survives, such that Eq. (8) restricted to that subset returns a vector $\vec{x}^{\star}$ with only positive entries. As first anticipated by the late Lord Robert May, not all ecologies can be stable [9].
@@
has a diverging eigenvalue $\lambda^{\star}$ in the large $N$ limit [55]. This means that any small change in the fitness of one species can have dramatic consequences on the whole system – in the present case, mass extinctions [10]. Similarly, the stability matrix $\mathbb{K}$, given by (cf. Eq. (3))
@@ 3.4 Stabilizing an unstable equilibrium
In other words, the better one is able to stabilize the system, the more difficult it becomes to predict its future evolution! This in turn generates occasional large errors, leading to large swings that are later corrected as the control policy becomes efficient again [64]. This rather beautiful scenario is quite universal – for example heart-beat regulation is a result of two conflicting forces: excitation vs. inhibition (sympathetic vs. parasympathetic nervous system), or brain activity regulation [68]. Such ideas are probably also relevant in economics and finance (see section 5.2, 5.4 below). In fact, the title of the present subsection is tantalizingly close to that of Minsky’s seminal book, “Stabilizing an Unstable Economy” [69].
@@ 4 Self-Organized Criticality in Economics
The aim of the present section is to revisit the issue and discuss more recent proposals that lead to self-organized criticality in economic systems, based on arguably less artificial assumptions. The hope is that putting the SOC scenario on firmer theoretical ground could revive the interest of economists. It is indeed fair to say that despite the high profile of its authors, the 1993 paper of Bak et al. [1] did not reach the level of attention that (in my view) it deserved  [^4] At the time of writing, the joint number of citations of Bak et al. 1993 [1] and Scheinkman-Woodford 1994 [21] is $\sim 800$ according to Google Scholar, compared for example to 2,050 for Gabaix’ 2011“granularity” paper [37], or 2,700 for Acemoglu, Carvalho et al. 2012 paper on the network origin of aggregate fluctuations [70]. – probably because of its purely conceptual nature, and the lack of precise, actionable proposal to fit on data.
@@ 4.2 The ecology of production networks (statics)
Expanding Eq. (11) to first order in $\zeta$, one notices that since $\sum_{j>0}a_{ij}<1$, the Perron-Frobenius theorem ensures that $q\to\infty$ Cobb-Douglas networked economies (such as those considered in [78, 70]) always have a feasible equilibrium where all firms can survive, for any network and any productivities. Therefore, the type of crisis that takes place for $q<+\infty$, where bankruptcies occur, has no counterpart in a Cobb-Douglas economy.